In Florida, when there were 447 boats registered (in thousands) there were 13 manatees killed that year. During a year when there were 719 boats registered (in thousands) there were 47 manatees killed that year. Assuming a linear relationship between boats registered and manatees killed, find a linear equation that describes the number of manatees killed with respect to boats registered. (discuss the meaning of the slope for this problem).
Answer by stanbon(48574)

In Florida, when there were 447 boats registered (in thousands) there were 13 manatees killed that year. During a year when there were 719 boats registered (in thousands) there were 47 manatees killed that year. Assuming a linear relationship between boats registered and manatees killed, find a linear equation that describes the number of manatees killed with respect to boats registered. (discuss the meaning of the slope for this problem).
-----------------------------
(447,13) and (719,47)
---------------------
slope: (47-13)/(719-447) = 34/272 = 0.125 = 1/8
==================
intercept: 13 = (1/8)447 + b
b = -42.875
--------------------
EQUATION:
# killed = (1/8)(boats in thousands) - 42.875
========================
Meaning of the slope.
When number of boats increases by 1000, number of manatees killed increases
by 1/8
=============
Cheers,
Stan H.